Please help me simplify this expression.
(3-2√7) (3+2√7)
FYI- The symbol “ √” is a radical expression (square root). I truly apprecaite any help you can give me on this.
Thanks !!!

Answer by stanbon(75887)

Please help me simplify this expression.
(3-2√7) (3+2√7)
---
Since the form is (a-b)(a+b) which has a product of a^2-b^2,
you have (3^1 - (2sqrt(7))^2) = 9 - 4*7 = 9-28 = -19
--------------------------------
Cheers,
Stan H.
=================================